Gasoline is C8H18. Calculate how much water and carbon dioxide is produced from 1 gallon of gasoline (SG=0.75).
1
Expert's answer
2012-03-01T10:05:40-0500
1 gallon = 3.785 L
M(C8H18)=114 g/mol
m(C8H18)=0.75 g/mL*3.785 L=2.84 g
n(C8H18)=m(C8H18)/M(C8H18)=2.84/114=0.025 mol
C8H18 + 12.5O2 = 8CO2 + 9H2O
n (CO2)=8*n(C8H18)=0.025 mol*8=0.2 mol
m(CO2)=0.2 mol*44g/mol=8.8 g
n(H2O)=9*n(C8H18)=0.025*9=0.225 mol
m(H2O)= 0.225 mol*18g/mol=4.05 g
Answer:
m(CO2)=8.8 g, m(H2O)=4.05 g.